# TabImpute: Accurate and Fast Zero-Shot Missing-Data Imputation with a Pre-Trained Transformer

### v3

@@ A.3 Additional tables
Pattern OT K-Nearest Neighbors MICE GAIN SoftImpute
MCAR 0.401 ± 0.034 0.323 ± 0.036 0.302 ± 0.036 0.318 ± 0.031 0.265 ± 0.041
NN-MNAR 0.334 ± 0.033 0.278 ± 0.031 0.254 ± 0.033 0.265 ± 0.026 0.223 ± 0.043
Self-Masking-MNAR 0.211 ± 0.034 0.207 ± 0.033 0.264 ± 0.044 0.224 ± 0.029 0.166 ± 0.030
Col-MAR 0.221 ± 0.031 0.227 ± 0.032 0.182 ± 0.030 0.156 ± 0.027 0.167 ± 0.033
Block-MNAR 0.199 ± 0.029 0.186 ± 0.029 0.155 ± 0.027 0.098 ± 0.016 0.165 ± 0.035
Seq-MNAR 0.192 ± 0.023 0.213 ± 0.027 0.139 ± 0.023 0.099 ± 0.015 0.181 ± 0.032
Panel-MNAR 0.247 ± 0.025 0.277 ± 0.028 0.146 ± 0.021 0.167 ± 0.023 0.181 ± 0.032
Polarization-MNAR 0.177 ± 0.018 0.146 ± 0.018 0.109 ± 0.015 0.109 ± 0.013 0.118 ± 0.021
Soft-Polarization-MNAR 0.108 ± 0.020 0.075 ± 0.018 0.066 ± 0.016 0.078 ± 0.014 0.077 ± 0.021
Latent-Factor-MNAR 0.306 ± 0.031 0.246 ± 0.031 0.227 ± 0.032 0.202 ± 0.023 0.194 ± 0.039
Cluster-MNAR 0.314 ± 0.033 0.254 ± 0.032 0.225 ± 0.032 0.239 ± 0.029 0.187 ± 0.038
Two-Phase-MNAR 0.260 ± 0.030 0.293 ± 0.031 0.191 ± 0.030 0.198 ± 0.026 0.190 ± 0.035
Censoring-MNAR 0.084 ± 0.016 0.095 ± 0.011 0.131 ± 0.022 0.127 ± 0.020 0.081 ± 0.021
Overall 0.235 ± 0.009 0.217 ± 0.008 0.184 ± 0.008 0.175 ± 0.007 0.169 ± 0.009
Pattern ForestDiffusion TabPFN MIWAE Col Mean
MCAR 0.220 ± 0.029 0.070 ± 0.008 0.036 ± 0.005 0.000 ± 0.000
NN-MNAR 0.177 ± 0.026 0.042 ± 0.006 0.025 ± 0.003 0.000 ± 0.000
Self-Masking-MNAR 0.116 ± 0.024 0.106 ± 0.016 0.026 ± 0.004 0.000 ± 0.000
Col-MAR 0.108 ± 0.020 0.081 ± 0.011 0.030 ± 0.003 0.000 ± 0.000
Block-MNAR 0.106 ± 0.017 0.051 ± 0.008 0.023 ± 0.003 0.000 ± 0.000
Seq-MNAR 0.089 ± 0.014 0.034 ± 0.003 0.022 ± 0.003 0.000 ± 0.000
Panel-MNAR 0.120 ± 0.016 0.083 ± 0.011 0.039 ± 0.006 0.000 ± 0.000
Polarization-MNAR 0.091 ± 0.012 0.036 ± 0.004 0.029 ± 0.004 0.000 ± 0.000
Soft-Polarization-MNAR 0.031 ± 0.004 0.004 ± 0.001 0.014 ± 0.001 0.000 ± 0.000
Latent-Factor-MNAR 0.152 ± 0.022 0.053 ± 0.007 0.028 ± 0.003 0.000 ± 0.000
Cluster-MNAR 0.181 ± 0.027 0.058 ± 0.009 0.033 ± 0.005 0.000 ± 0.000
Two-Phase-MNAR 0.126 ± 0.022 0.086 ± 0.012 0.022 ± 0.003 0.000 ± 0.000
Censoring-MNAR 0.071 ± 0.011 0.072 ± 0.018 0.067 ± 0.018 0.000 ± 0.000
Overall 0.122 ± 0.006 0.060 ± 0.003 0.030 ± 0.002 0.000 ± 0.000
Table 9: Mean Normalized Negative RMSE ± Standard Error by Missingness Pattern. Note that these normalized values are slightly different than Tab. 1 because the nonlinear model is included here.
@@
Self-Masking-MNAR Coefficient set $\alpha$ $\{-2,-1,1,2\}$
@@
Table 12: Non-normalized RMSE values for MCAR pattern by dataset. Dataset columns are standardized based on observed values (mean 0, variance 1).
Dataset TabImpute EWF-TabPFN HyperImpute MissForest
EgyptianSkulls 0.901 0.917 0.961 0.946
FacultySalaries 0.607 0.693 0.811 0.567
Glass-Classification 0.889 0.937 0.810 0.886
ICU 1.059 1.045 1.137 1.112
MercuryinBass 1.311 1.373 1.377 1.343
Rainfall-in-Kerala-1901-2017 0.908 0.944 0.902 0.950
SMSA 0.840 0.814 0.939 0.786
SolarPower 0.889 0.805 0.873 0.888
Student-Scores 0.429 0.392 0.441 0.435
Swiss-banknote-conterfeit-detection 0.774 0.976 0.856 0.730
WineDataset 0.829 0.928 0.803 0.797
alcohol-qcm-sensor 0.467 0.590 0.580 0.517
analcatdata_election2000 0.364 0.538 0.662 0.615
analcatdata_gviolence 0.687 0.796 0.801 0.877
analcatdata_olympic2000 1.950 1.871 1.922 1.895
appendicitis_test 0.694 0.861 0.650 0.792
appendicitis_test_edsa 0.534 0.668 0.532 0.569
autoPrice 0.735 0.834 0.762 0.702
baskball 1.008 1.028 1.100 1.109
benzo32 1.070 1.055 1.055 1.053
breast-cancer-coimbra 1.068 1.058 1.183 1.098
chatfield_4 0.473 0.476 0.531 0.492
chscase_vine1 0.928 0.809 0.996 0.833
dataset_analcatdata_creditscore 1.238 1.063 1.131 1.122
divorce_prediction 0.542 0.525 0.528 0.467
edm 0.612 0.763 0.553 0.566
humans_numeric 0.989 0.952 0.970 0.991
machine_cpu 0.902 1.039 0.956 0.858
metafeatures 2.609 2.263 2.090 2.225
pollution 1.946 1.967 2.065 1.837
pwLinear 1.031 1.075 1.356 1.356
pyrim 0.847 0.877 0.863 0.859
rabe_131 1.016 0.792 1.003 1.031
rabe_148 0.905 0.985 1.017 0.958
rabe_265 1.104 1.131 1.298 1.090
sleuth_case1201 0.940 1.054 0.939 0.963
sleuth_ex1605 1.026 1.087 1.257 1.071
slump 0.750 0.871 0.650 0.805
treepipit 1.226 1.265 1.169 1.172
visualizing_hamster 0.754 0.872 0.905 0.750
wisconsin 0.854 0.777 0.685 0.783
witmer_census_1980 0.654 0.704 0.616 0.637
Table 13: Non-normalized RMSE Values for MCAR Pattern by Dataset including ReMasker. Note that we could not replicate ReMasker’s results using their open-source implementation. In order to not misrepresent their method, we do not include this table in our main results.
Dataset TabImpute HyperImpute MissForest ReMasker
EgyptianSkulls 0.901 0.961 0.946 2.329
FacultySalaries 0.607 0.811 0.567 1.534
Glass-Classification 0.889 0.810 0.886 3.126
ICU 1.059 1.137 1.112 1.479
MercuryinBass 1.311 1.377 1.343 2.198
Rainfall-in-Kerala-1901-2017 0.908 0.902 0.950 3.560
SMSA 0.840 0.939 0.786 1.979
SolarPower 0.889 0.873 0.888 1.589
Student-Scores 0.429 0.441 0.435 1.785
Swiss-banknote-conterfeit-detection 0.774 0.856 0.730 1.364
WineDataset 0.829 0.803 0.797 1.454
alcohol-qcm-sensor 0.467 0.580 0.517 2.044
analcatdata_election2000 0.364 0.662 0.615 4.563
analcatdata_gviolence 0.687 0.801 0.877 1.617
analcatdata_olympic2000 1.950 1.922 1.895 2.787
appendicitis_test 0.694 0.650 0.792 2.174
appendicitis_test_edsa 0.534 0.532 0.569 1.986
autoPrice 0.735 0.762 0.702 2.190
baskball 1.008 1.100 1.109 2.246
benzo32 1.070 1.055 1.053 3.282
breast-cancer-coimbra 1.068 1.183 1.098 1.820
chatfield_4 0.473 0.531 0.492 3.545
chscase_vine1 0.928 0.996 0.833 2.720
dataset_analcatdata_creditscore 1.238 1.131 1.122 1.469
divorce_prediction 0.542 0.528 0.467 1.216
edm 0.612 0.553 0.566 1.415
humans_numeric 0.989 0.970 0.991 4.165
machine_cpu 0.902 0.956 0.858 1.402
metafeatures 2.609 2.090 2.225 2.715
pollution 1.946 2.065 1.837 2.985
pwLinear 1.031 1.356 1.356 1.040
pyrim 0.847 0.863 0.859 1.647
rabe_131 1.016 1.003 1.031 2.076
rabe_148 0.905 1.017 0.958 2.508
rabe_265 1.104 1.298 1.090 2.633
sleuth_case1201 0.940 0.939 0.963 2.455
sleuth_ex1605 1.026 1.257 1.071 2.327
slump 0.750 0.650 0.805 1.960
treepipit 1.226 1.169 1.172 1.648
visualizing_hamster 0.754 0.905 0.750 1.884
wisconsin 0.854 0.685 0.783 2.422
witmer_census_1980 0.654 0.616 0.637 2.487

### v4

@@ A.4 Additional tables
Pattern OT CACTI K-NN MICE GAIN
MCAR 0.401 ± 0.034 0.337 ± 0.037 0.323 ± 0.036 0.302 ± 0.036 0.318 ± 0.031
NN-MNAR 0.334 ± 0.033 0.263 ± 0.036 0.278 ± 0.031 0.254 ± 0.033 0.265 ± 0.026
Self-MNAR 0.211 ± 0.034 0.203 ± 0.032 0.207 ± 0.033 0.264 ± 0.044 0.224 ± 0.029
Col-MAR 0.221 ± 0.031 0.242 ± 0.030 0.227 ± 0.032 0.182 ± 0.030 0.156 ± 0.027
Block-MNAR 0.199 ± 0.029 0.179 ± 0.031 0.186 ± 0.029 0.155 ± 0.027 0.098 ± 0.016
Seq-MNAR 0.192 ± 0.023 0.186 ± 0.032 0.213 ± 0.027 0.139 ± 0.023 0.099 ± 0.015
Panel-MNAR 0.247 ± 0.025 0.243 ± 0.029 0.277 ± 0.028 0.146 ± 0.021 0.167 ± 0.023
Polarization-MNAR 0.177 ± 0.018 0.150 ± 0.019 0.146 ± 0.018 0.109 ± 0.015 0.109 ± 0.013
Soft-Polarization-MNAR 0.108 ± 0.020 0.114 ± 0.022 0.075 ± 0.018 0.066 ± 0.016 0.078 ± 0.014
Latent-Factor-MNAR 0.306 ± 0.031 0.253 ± 0.033 0.246 ± 0.031 0.227 ± 0.032 0.202 ± 0.023
Cluster-MNAR 0.314 ± 0.033 0.280 ± 0.035 0.254 ± 0.032 0.225 ± 0.032 0.239 ± 0.029
Two-Phase-MNAR 0.260 ± 0.030 0.285 ± 0.036 0.293 ± 0.031 0.191 ± 0.030 0.198 ± 0.026
Censoring-MNAR 0.084 ± 0.016 0.103 ± 0.012 0.095 ± 0.011 0.131 ± 0.022 0.127 ± 0.020
Overall 0.235 ± 0.009 0.218 ± 0.009 0.217 ± 0.008 0.184 ± 0.008 0.175 ± 0.007
Pattern ReMasker SoftImpute ForestDiffusion Col-TabPFN MIWAE
MCAR 0.327 ± 0.036 0.265 ± 0.041 0.220 ± 0.029 0.070 ± 0.008 0.036 ± 0.005
NN-MNAR 0.196 ± 0.032 0.223 ± 0.043 0.177 ± 0.026 0.042 ± 0.006 0.025 ± 0.003
Self-MNAR 0.151 ± 0.026 0.166 ± 0.030 0.116 ± 0.024 0.106 ± 0.016 0.026 ± 0.004
Col-MAR 0.185 ± 0.032 0.167 ± 0.033 0.108 ± 0.020 0.081 ± 0.011 0.030 ± 0.003
Block-MNAR 0.149 ± 0.031 0.165 ± 0.035 0.106 ± 0.017 0.051 ± 0.008 0.023 ± 0.003
Seq-MNAR 0.111 ± 0.021 0.181 ± 0.032 0.089 ± 0.014 0.034 ± 0.003 0.022 ± 0.003
Panel-MNAR 0.186 ± 0.027 0.181 ± 0.032 0.120 ± 0.016 0.083 ± 0.011 0.039 ± 0.006
Polarization-MNAR 0.119 ± 0.019 0.118 ± 0.021 0.091 ± 0.012 0.036 ± 0.004 0.029 ± 0.004
Soft-Polarization-MNAR 0.046 ± 0.012 0.077 ± 0.021 0.031 ± 0.004 0.004 ± 0.001 0.014 ± 0.001
Latent-Factor-MNAR 0.177 ± 0.029 0.194 ± 0.039 0.152 ± 0.022 0.053 ± 0.007 0.028 ± 0.003
Cluster-MNAR 0.212 ± 0.034 0.187 ± 0.038 0.181 ± 0.027 0.058 ± 0.009 0.033 ± 0.005
Two-Phase-MNAR 0.263 ± 0.038 0.190 ± 0.035 0.126 ± 0.022 0.086 ± 0.012 0.022 ± 0.003
Censoring-MNAR 0.080 ± 0.012 0.081 ± 0.021 0.071 ± 0.011 0.072 ± 0.018 0.067 ± 0.018
Overall 0.169 ± 0.008 0.169 ± 0.009 0.122 ± 0.006 0.060 ± 0.003 0.030 ± 0.002
Table 9: Mean Normalized Negative RMSE ± Standard Error by Missingness Pattern including a nonlinear factor model-based TabImpute. Note that these normalized values are slightly different than Tab. 1 because the nonlinear model is included here.
@@
Self-MNAR Coefficient set $\alpha$ $\{-2,-1,1,2\}$
@@
Table 12: Non-normalized RMSE Values for MCAR (Probability of Missing = 0.4) by Dataset
Dataset TabImpute EWF-TabPFN HyperImpute MissForest OT
EgyptianSkulls 0.995 1.019 1.057 1.047 1.074
FacultySalaries 0.605 0.795 0.720 0.615 0.638
Glass-Classification 0.865 0.900 0.834 0.864 0.842
ICU 1.006 0.992 1.080 1.057 0.993
MercuryinBass 0.768 0.870 0.846 0.791 0.808
Rainfall-in-Kerala-1901-2017 0.876 0.918 0.855 0.929 0.924
SMSA 0.941 0.911 1.069 0.893 0.935
SolarPower 0.876 0.785 0.862 0.878 0.907
Student-Scores 0.431 0.441 0.425 0.471 0.520
Swiss-banknote-conterfeit-detection 0.731 0.926 0.805 0.664 0.732
WineDataset 0.760 0.866 0.738 0.732 0.723
alcohol-qcm-sensor 0.358 0.449 0.408 0.392 0.504
analcatdata_election2000 0.738 2.969 3.454 5.660 5.869
analcatdata_gviolence 0.757 0.812 0.873 0.940 0.826
analcatdata_olympic2000 0.622 0.550 0.600 0.587 0.617
appendicitis_test 0.631 0.787 0.588 0.726 0.692
appendicitis_test_edsa 0.591 0.735 0.610 0.660 0.653
autoPrice 0.631 0.735 0.657 0.618 0.646
baskball 0.969 0.995 1.063 1.074 0.987
benzo32 1.019 1.003 1.001 1.000 0.889
breast-cancer-coimbra 0.937 0.913 1.068 0.964 0.939
chatfield_4 0.421 0.480 0.480 0.444 0.476
chscase_vine1 0.809 0.692 0.885 0.746 0.841
dataset_analcatdata_creditscore 1.172 1.006 1.084 1.069 1.006
divorce_prediction 0.510 0.480 0.484 0.433 0.458
edm 0.602 0.767 0.536 0.555 0.561
humans_numeric 0.966 1.004 0.985 0.994 1.014
machine_cpu 0.779 0.927 0.824 0.729 0.818
metafeatures 1.437 2.471 3.294 4.329 3.578
pollution 0.941 0.905 1.087 0.871 0.911
pwLinear 0.999 1.038 1.310 1.311 1.058
pyrim 0.732 0.802 0.666 0.678 0.761
rabe_131 0.966 0.636 0.944 0.986 0.926
rabe_148 0.828 0.856 1.232 0.953 0.984
rabe_265 1.033 1.060 1.169 1.017 1.007
sleuth_case1201 0.834 0.943 0.862 0.891 0.895
sleuth_ex1605 0.973 1.045 1.144 1.026 0.932
slump 0.835 0.954 0.737 0.894 0.874
treepipit 0.980 1.044 0.909 0.907 0.917
visualizing_hamster 0.817 0.921 1.019 0.821 0.849
wisconsin 0.731 0.645 0.555 0.661 0.702
witmer_census_1980 0.906 0.988 0.864 0.894 0.898